OpenStudy (anonymous):

can somebody tell me if i have 3 i need to multiple it by 100 wht number can i multiple it by

OpenStudy (pdoctor):

What?

OpenStudy (jamesj):

You want to know this? \[3 \times 100 \]

OpenStudy (anonymous):

so u said i can multiply 3*1 and it will get me 100

OpenStudy (jamesj):

No, that's not right. What is your question? What you wrote originally doesn't make any sense.

OpenStudy (anonymous):

well im trying to convert a fraction to a percent

OpenStudy (anonymous):

and the fraction is 5/3

OpenStudy (jamesj):

Ok. So in percent this is \[\frac{5}{3} \times 100 = \frac{500}{3} = 166 \frac{2}{3}\]

OpenStudy (anonymous):

OK BUT IM DOING AND SHORT EXTENDED RESPONSE

OpenStudy (anonymous):

ARE U THERE U CHANGE IT INTO A FRACTION I NEED IT TO HAVE A PERCENT CAN U TELL ME IT

OpenStudy (anonymous):

THE PERCENT

hero (hero):

\[3\dot\ \frac{100}{3} = 100\]

OpenStudy (anonymous):

no tht give u 3 |

OpenStudy (anonymous):

no this is wht i go so far

OpenStudy (anonymous):

5/3=100/?

OpenStudy (jamesj):

Oh, completely different question. You're asking: What is the x such that 5/3 = 100/x ? Well, start 5/3 and multiply top and bottom by 20:\[\frac{5}{3} = \frac{5 \times 20}{3 \times 20} = \frac{100}{60}\]

OpenStudy (anonymous):

ok so tht the percent #

OpenStudy (anonymous):

60%

OpenStudy (anonymous):

is yell there\